OpenStudy (mortonsalt):

A satellite is placed in orbit around the Sun. The orbital radius of the satellite is twice the orbital radius of the Earth. What is the orbital radius of this satellite?

OpenStudy (mortonsalt):

Can I use \[\frac{ GMm }{ r^2 } = \frac {m4\pi^2r}{T^2}\]

OpenStudy (anonymous):

The orbital radius of the Earth around the sun is 1 AU. Twice that radius is 2 AU.